How many liters of a 24 percent alcohol solution must be mixed with 20 liters of a 46 percent alcohol solution to get a 28 percent alcohol solution?

Answer by ankor@dixie-net.com(22740)

How many liters of a 24 percent alcohol solution must be mixed with 20 liters of a 46 percent alcohol solution to get a 28 percent alcohol solution?
:
let a = amt of 24% alcohol required
:
.24a + .46(20) = .28(a+20)
.24a + 9.2 = .28a + 5.6
9.2 - 5.6 = .28a - .24a
3.6 = .04a
a = 3.6/.04
a = 90 liters of 24% alcohol required
:
Check
.24(90) + .46(20) = .28(90+20)
21.6 + 9.2 = .28(110)
30.2 = 30.2